# 0.5 eV QCD Axion Dark Matter

### v41

@@ 2.1 Chirality and Helicity: What’s the Difference?
One can easily imagine a right-handed matter field of flavor $M$ that feels the weak nuclear force. In our notation, such a field is denoted $M_{R}^{-}$. Its antiparticle in the same notation is $\overline{M}_{L}^{-}$. This means the antiparticle has the same flavor $M$ and chirality $P=-$ but opposite helicity $H=L$. The chirality $P=-$ indicates that both particle $M_{R}^{-}$ and the antiparticle $\overline{M}_{L}^{-}$ feel the weak nuclear force. The flip in helicity means the antiparticle is left-handed when the particle is right-handed.
@@
To make the model of 0.5 eV QCD axion dark matter, then, we have gone beyond the standard model of particle physics. We include “normal” left-handed matter $M_{L}^{-}$ that feels the weak nuclear force and “normal” right-handed matter $M_{R}^{+}$ that does not feel it along with their antiparticles. Similarly included are right-handed “normal” antimatter $\overline{M}_{R}^{-}$ which feels the weak nuclear force and left-handed “normal” antimatter $\overline{M}_{L}^{+}$ which does not feel it. But also, we allow “mirror” matter ($M_{L}^{+}$ and $M_{R}^{-}$) as well as “mirror” antimatter ($\overline{M}_{R}^{+}$ and $\overline{M}_{L}^{-}$).
We note that mirror matter has the same correlation between helicity $H=L,R$ and chirality $P=\pm$ as normal antimatter. Left-handed mirror matter $M_{L}^{+}$ and left-handed normal antimatter $\overline{M}_{L}^{+}$ both do not feel the weak nuclear force, while right-handed mirror matter $M_{R}^{-}$ and right-handed normal antimatter $\overline{M}_{R}^{-}$ both feel it. Similarly, normal matter has the same correlation between helicity and chirality as mirror antimatter. Left-handed normal matter $M_{L}^{-}$ and left-handed mirror antimatter $\overline{M}_{L}^{-}$ both feel the weak nuclear force, while right-handed normal matter $M_{R}^{+}$ and right-handed mirror antimatter $\overline{M}_{R}^{+}$ both do not feel it.
@@ 2.3 Discrete Space-Time Symmetries: $\mathcal{P}$, $\mathcal{C}$, $\mathcal{T}$ and All That
$\mathcal{C}M_{H}^{P}$ $=$ $\overline{M}_{H}^{\overline{P}}~{}~{}~{}\mathcal{C}\overline{M}_{H}^{P}=M_{H}^{\overline{P}}$
$\mathcal{T}M_{H}^{P}$ $=$ $\overline{M}_{\overline{H}}^{P}~{}~{}~{}\mathcal{T}\overline{M}_{H}^{P}=M_{\overline{H}}^{P},$ (2.3)
@@
Three key features of the discrete space-time symmetries determine the form of the transformation given in Eq. (2.3). First, space-inversion flips helicity $H$ since linear momentum is even under space-inversion but spin is odd under $\mathcal{P}$. However, the weak nuclear force breaks space-inversion $\mathcal{P}$ [28, 29]. Given the correlation between helicity $H$ and chirality $P$ for both normal and mirror fields, this means that chirality $P$ must also flip under space-inversion to insure that matter which feels the weak nuclear force is interchanged with matter that does not feel this force. Since space-inversion takes matter into matter and antimatter into anti-matter, these features suffice to determine the form of $\mathcal{P}$ given in Eq. (2.3).
Second, charge-conjugation interchanges matter with antimatter and it is also broken by the weak interaction [28, 29]. In our notation, it follows that the chirality $P$ must flip under $\mathcal{C}$ to take fields that feel the weak nuclear force into fields that do not feel it. However, helicity $H$ is unchanged. Combining these features, we obtain the $\mathcal{C}$ transformation law shown in Eq. (2.3).
Third, the product $\mathcal{P}\mathcal{C}\mathcal{T}$ of all three symmetries is the identity in the sense that $\mathcal{P}\mathcal{C}\mathcal{T}M_{H}^{P}=M_{H}^{P}$ and $\mathcal{P}\mathcal{C}\mathcal{T}\overline{M}_{H}^{P}=\overline{M}_{H}^{P}$. This relation is respected by all three gauge couplings in the standard model and has been observed to hold experimentally to high precision for a wide variety of particle properties including charge, mass, life-time, magnetic dipole moment, and so on, for a wide range of particles [42]. The transformation law for $\mathcal{T}$ in Eq. (2.3) is then fixed by simply requiring $\mathcal{T}=\mathcal{P}\mathcal{C}$ and using the laws for $\mathcal{P}$ and $\mathcal{C}$ given in Eq. (2.3). We find $\mathcal{T}$ reverses helicity $H$, exchanges matter for antimatter, and leaves unchanged chirality $P$.
@@ 3.2 Big Bang Temperature from QCD Axion Angle Jump $\Delta\theta_{A}=\pi$
The factor of six in the denominator on the left-hand side of Eq. (3.4) is simply the number of flavors of quarks. This factor follows from the conventional normalization of the axion-gluon coupling together with the form of the continuous $U(1)_{A}$ axial phase symmetry given in Eq. (LABEL:compact_phase) [27, 48]. The symmetry acts in the same way on all flavors $M$ of quarks and this leads to the factor of six in the denominator on the right-hand side of Eq. (3.4).
@@
where $\tau_{\pi}=(8.43\pm 0.13)\times 10^{-17}{\rm sec}/(0.98823)=(8.53\pm 0.13)\times 10^{-17}~{}{\rm nsec}$ [23] is the two-photon lifetime of the neutral pion $\pi^{0}$, $f_{\pi}=(1/\sqrt{2})(127.13/0.974)=92.3~{}{\rm MeV}$ [23] is the pion decay constant, and $m_{\pi}=134.98~{}{\rm MeV}$ [23] is the rest-mass energy of the neutral pion $\pi^{0}$.
@@ 6.4 Comment on White Dwarfs, Red Giants, Pulsars, and Supernova 1987A
The strength of the coupling $g_{Aee}$ of 0.5 eV QCD axion dark matter to electrons appears to be precisely zero $g_{Aee}=0$ based on analysis of the unobserved effects that such an axion-electron coupling would have on the cooling of white dwarfs and red giants were the strength $g_{Aee}$ not precisely zero [27]. By contrast, the cooling of pulsars and the energy flow in the collapsing core of the progenitor star for supernova 1987 A appear to both be dominated by the effects of the non-zero axion-nucleon coupling to 0.5 eV QCD axion dark matter [27]. Indeed, the strength of the axion-nucleon coupling $g_{ANN}f_{A}\approx g_{\pi NN}f_{\pi}$ [48] is roughly set by the axion decay constant $f_{A}$ in Eq. (3.2) and by the product $g_{\pi NN}f_{\pi}\approx g_{A}m_{N}c^{2}$ [23] of the pion-nucleon coupling strength $g_{\pi NN}$ with the pion decay constant $f_{\pi}$, where $g_{A}=-1.2754\pm 0.0013$ [23] is the axial weak nuclear coupling strength and $m_{N}\approx 940~{}{\rm MeV/c^{2}}$ is the nucleon rest-mass.

### v42

@@ 2.1 Chirality and Helicity: What’s the Difference?
One can easily imagine a right-handed matter field of flavor $M$ that feels the weak nuclear force. In our notation, such a field is denoted $M_{R}^{-}$. Its antiparticle in the same notation is $\overline{M}_{R}^{-}$. This means the antiparticle has the same flavor $M$ and chirality $P=-$ but opposite helicity. The chirality $P=-$ indicates that both particle $M_{R}^{-}$ and the antiparticle $\overline{M}_{R}^{-}$ feel the weak nuclear force. The flip in helicity means the antiparticle is left-handed when the particle is right-handed.
@@
To make the model of 0.5 eV QCD axion dark matter, then, we have gone beyond the standard model of particle physics. We include “normal” left-handed matter $M_{L}^{-}$ that feels the weak nuclear force and “normal” right-handed matter $M_{R}^{+}$ that does not feel it along with their antiparticles. Similarly included are right-handed “normal” antimatter $\overline{M}_{L}^{-}$ which feels the weak nuclear force and left-handed “normal” antimatter $\overline{M}_{R}^{+}$ which does not feel it. But also, we allow “mirror” matter ($M_{L}^{+}$ and $M_{R}^{-}$) as well as “mirror” antimatter ($\overline{M}_{R}^{+}$ and $\overline{M}_{L}^{-}$).
We note that mirror matter has the same correlation between helicity $H=L,R$ and chirality $P=\pm$ as normal antimatter. Left-handed mirror matter $M_{L}^{+}$ and left-handed normal antimatter $\overline{M}_{R}^{+}$ both do not feel the weak nuclear force, while right-handed mirror matter $M_{R}^{-}$ and right-handed normal antimatter $\overline{M}_{L}^{-}$ both feel it. Similarly, normal matter has the same correlation between helicity and chirality as mirror antimatter. Left-handed normal matter $M_{L}^{-}$ and left-handed mirror antimatter $\overline{M}_{R}^{-}$ both feel the weak nuclear force, while right-handed normal matter $M_{R}^{+}$ and right-handed mirror antimatter $\overline{M}_{L}^{+}$ both do not feel it.
@@ 2.3 Discrete Space-Time Symmetries: $\mathcal{P}$, $\mathcal{C}$, $\mathcal{T}$ and All That
$\mathcal{C}M_{H}^{P}$ $=$ $\overline{M}_{\overline{H}}^{\overline{P}}~{}~{}~{}\mathcal{C}\overline{M}_{H}^{P}=M_{\overline{H}}^{\overline{P}}$
$\mathcal{T}M_{H}^{P}$ $=$ $\overline{M}_{H}^{P}~{}~{}~{}\mathcal{T}\overline{M}_{H}^{P}=M_{H}^{P},$ (2.3)
@@
Three key features of the discrete space-time symmetries determine the form of the transformation given in Eq. (2.3). First, space-inversion flips helicity $H$ since linear momentum is odd under space-inversion but spin is even under $\mathcal{P}$. However, the weak nuclear force breaks space-inversion $\mathcal{P}$ [28, 29]. Given the correlation between helicity $H$ and chirality $P$ for both normal and mirror fields, this means that chirality $P$ must also flip under space-inversion to insure that matter which feels the weak nuclear force is interchanged with matter that does not feel this force. Since space-inversion takes matter into matter and antimatter into anti-matter, these features suffice to determine the form of $\mathcal{P}$ given in Eq. (2.3).
Second, charge-conjugation interchanges matter with antimatter and it is also broken by the weak interaction [28, 29]. In our notation, it follows that the chirality $P$ must flip under $\mathcal{C}$ to take fields that feel the weak nuclear force into fields that do not feel it. However, helicity is unchanged which means the label $H$ must flip in our notation. Combining these features, we obtain the $\mathcal{C}$ transformation law shown in Eq. (2.3).
Third, the product $\mathcal{P}\mathcal{C}\mathcal{T}$ of all three symmetries is the identity in the sense that $\mathcal{P}\mathcal{C}\mathcal{T}M_{H}^{P}=M_{H}^{P}$ and $\mathcal{P}\mathcal{C}\mathcal{T}\overline{M}_{H}^{P}=\overline{M}_{H}^{P}$. This relation is respected by all three gauge couplings in the standard model and has been observed to hold experimentally to high precision for a wide variety of particle properties including charge, mass, life-time, magnetic dipole moment, and so on, for a wide range of particles [42]. The transformation law for $\mathcal{T}$ in Eq. (2.3) is then fixed by simply requiring $\mathcal{T}=\mathcal{P}\mathcal{C}$ and using the laws for $\mathcal{P}$ and $\mathcal{C}$ given in Eq. (2.3). We find $\mathcal{T}$ reverses helicity, exchanges matter for antimatter, and leaves unchanged chirality.
@@ 3.2 Big Bang Temperature from QCD Axion Angle Jump $\Delta\theta_{A}=\pi$
The factor of six in the denominator on the left-hand side of Eq. (3.4) is simply the number of flavors of quarks. This factor follows from the conventional normalization of the axion-gluon coupling together with the form of the continuous $U(1)_{A}$ axial phase symmetry given in Eq. (2.2) [27, 48]. The symmetry acts in the same way on all flavors $M$ of quarks and this leads to the factor of six in the denominator on the right-hand side of Eq. (3.4).
@@
where $\tau_{\pi}=(8.43\pm 0.13)\times 10^{-17}{\rm sec}/(0.98823)=(8.53\pm 0.13)\times 10^{-17}~{}{\rm sec}$ [23] is the two-photon lifetime of the neutral pion $\pi^{0}$, $f_{\pi}=(1/\sqrt{2})(127.13/0.974)=92.3~{}{\rm MeV}$ [23] is the pion decay constant, and $m_{\pi}=134.98~{}{\rm MeV}$ [23] is the rest-mass energy of the neutral pion $\pi^{0}$.
@@ 6.4 Comment on White Dwarfs, Red Giants, Pulsars, and Supernova 1987A
The brightness of the tip of the red-giant branch in globular clusters and the white dwarf luminosity function both indicate that the “bare” or “tree-level” coupling of electrons to 0.5 eV QCD axion dark matter is precisely zero and that the leading order axion-electron coupling comes in only at “loop-level” [27]. By contrast, the cooling of pulsars and the energy flow in the collapsing core of the progenitor star for supernova 1987 A appear to both be dominated by the effects of the non-zero axion-nucleon coupling to 0.5 eV QCD axion dark matter [27]. Indeed, the strength of the axion-nucleon coupling $g_{ANN}f_{A}\approx g_{\pi NN}f_{\pi}$ [48] is roughly set by the axion decay constant $f_{A}$ in Eq. (3.2) and by the product $g_{\pi NN}f_{\pi}\approx g_{A}m_{N}c^{2}$ [23] of the pion-nucleon coupling strength $g_{\pi NN}$ with the pion decay constant $f_{\pi}$, where $g_{A}=-1.2754\pm 0.0013$ [23] is the axial weak nuclear coupling strength and $m_{N}\approx 940~{}{\rm MeV/c^{2}}$ is the nucleon rest-mass.